# StereoDiffusion: Training-Free Stereo Image Generation Using Latent Diffusion Models

### v1

@@ 3.3 Self-Attention layers modification
As numerous studies have attempted to modify the attention mechanisms within Stable Diffusion to achieve the goal of modifying the original images [17, 6, 25, 19, 24], we tackled this challenge by utilizing both Unidirectional and Bidirectional Self-Attention mechanisms. This method eliminates the need to fine-tune the model to adjust its weights.
In the case of self-attention, non-rigid editing cannot be performed as the semantic layout and structures are maintained. Similar to sharing semantic information between different samples in the same batch using 3D convolution to align content across batches in video generation tasks [22, 23, 42], applying self-attention between samples within the same batch has a comparable effect [24, 35]. Querying the left-side image using the key and value of the right-side image in a unidirectional manner, enhancing the alignment from right image to left, is termed unidirectional self-attention. In contrast, employing queries from both the left and right sides to mutually query each other is referred to as bidirectional self-attention. However, bidirectional self-attention has a significant drawback: it aligns the left and right images with each other, thereby altering the input left-side image. Although this can enhance alignment, it is not a suitable option when users wish to keep the input image unchanged. Thus, despite its potential to improve alignment, the bidirectional approach may not be preferable if it is crucial to maintain the integrity of the input image.
The algorithm is shown in Algorithm 1. The term $\epsilon_{\theta}((z_{t},z^{\prime}_{t}),t,\mathcal{C})$ represents the computation of a single step $t$ of the diffusion process, which yields the noisy image $z_{t-1}$ and the attention map $M_{t}$. Here, $(z_{t},z^{\prime}_{t})$ denote the left and right latent variables respectively. In practical implementation, these latent variables are stacked together along the batch size dimension. However, they are represented separately here for ease of explanation. The expression $\epsilon_{\theta}((z_{t},{z^{\prime}_{t}}),t,\mathcal{C})\{M^{\prime}\leftarrow\widehat{M^{\prime}_{t}}\}$ denotes the diffusion step where the attention map $M$ is superseded by an additional given map $\widehat{M}$. We define the function $\operatorname{Edit}\left(M_{t},M^{\prime}_{t},t\right)$ as a general edit function, designed to process the $t$th attention maps of left and right latent variables.
@@
3.4 Application scenarios
As shown in Fig. 1, our method is compatible with various types of Stable Diffusion models, enabling it to: (a) produce the corresponding right-side image from an existing left-side image; (b) generate stereo images from text prompts; (c) produce the corresponding right-side image from an existing left-side image, where the pair shares the same composition but differs in content. For text-to-image and depth-to-image tasks, the initial noise is randomly generated. Thus, it is sufficient to apply a pixel shift to the denoised noise after a specific denoising step, as illustrated in Fig. 2. However, for generating stereo image pairs of an existing image, it is necessary to use null-text inversion [43] to obtain the latent space of the original image. A straightforward inversion technique was proposed for the DDIM sampling [44, 16]. This technique is grounded in the hypothesis that the ordinary differential equation (ODE) process is reversible, especially in scenarios involving small step sizes. The diffusion process is executed in reverse, meaning the transition is from $z_{0}$ to $z_{T}$, contrary to the typical $z_{T}$ to $z_{0}$ progression:
$z_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\,z_{t}+\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)\,\varepsilon_{\theta}\!\left(z_{t},t,\mathcal{C}\right).$ (11)
To address the inefficiency of mapping each noise vector to a single image, this method initiates with a default DDIM inversion at $w=1$ as its pivot trajectory. Subsequently, it optimizes around this trajectory using a standard guidance ratio of $w>1$. In practical applications, individual optimizations are conducted for each step $t$ during the diffusion process, aiming to closely approximate the initial trajectory $z^{*}$:
$\min\left\|z_{t-1}^{*}-z_{t-1}\right\|_{2}^{2}\,,$ (12)
where $z_{t-1}$ represents the intermediate result of the optimization. The approach involves substituting the default blank text embedding with an optimized embedding. This is due to a key characteristic of classifier-free guidance, which is significantly influenced by the unconditional prediction.
4 Experiments
We have compared our results with traditional methods such as ’leave blank’ and ’stretch’. Additionally, we have selected the 3D Photography techniques of Shih et al. [12] for comparison, as well as the RePaint method of Lugmayr et al. [45], which involves using Stable Diffusion for inpainting images processed by the traditional ’leave blank’ method. It is important to emphasize that RePaint is not inherently designed for generating stereo image pairs. However, we believe that employing inpainting techniques to fill in the blank areas after creating stereo images is a very straightforward and common approach. Thus, we have chosen to compare with the latest model that achieves good results in various metrics within the same Stable Diffusion framework. This comparison is intended to demonstrate the innovation and advantage of our method.
4.1 Quantitative evaluation
@@
Figure 5: Comparison of the same image generated using different methods. The rows present, respectively, the image with the lowest (worst) SSIM score, the image closest to the average SSIM score, and the image with the highest (best) SSIM score generated using our method. The other methods are represented solely by their results on specific images and do not necessarily reflect the best, average, or worst SSIM scores achievable by those methods. This approach is adopted to facilitate a direct comparison of the effects of each method on the same image. We also annotate the generated images with their LPIPS scores for reference. A ’Details Comparison’ is provided for a detailed comparison of images generated by the primary benchmark methods.
@@
Table 1: Quantitative evaluation results for Middlebury and KITTI: the results of generating right-side images from left-side images and disparity maps using different methods. We assess the similarity between the generated and the original images using PSNR, SSIM, and LPIPS. ’GT’ indicates the use of ground truth disparity maps, while ’pseudo’ denotes the use of disparity maps generated by a depth estimation model. Scores presented in bold indicate the best performance. The numbers in the top right represent the best scores, while those in the bottom right indicate the worst scores.
@@
PNSR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PNSR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
Since there is currently no metrics specifically for the stereo image pair generation, we quantitatively evaluate our results using the Middlebury [46] and KITTI [47] datasets. We evaluate the performance by generating the right-side image from the left-side image and its disparity map, and then comparing the model-generated right-side image with the ground truth image. We calculated the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) between the generated image and the ground truth. The results are in Table 1. We provide the settings used for each method in a supplemental document.
The use of null-text inversion [43] technique inherently causes distortion in images. On the Middlebury dataset, reference scores (for images generated by Stable Diffusion to be the same as the input) are: PSNR = 27.967, SSIM = 0.847, LPIPS = 0.046. The reference scores for the KITTI dataset are: PSNR = 25.615, SSIM = 0.762, LPIPS = 0.072. These scores represent the best possible outcomes achievable with the method we proposed. The quantitative analysis results, as seen in Table 1, indicate that our proposed method achieves state-of-the-art scores on both the datasets. Furthermore, as illustrated in Fig. 4, we selected images representing the best LPIPS, those closest to the average LPIPS, and the worst LPIPS from each method. This selection was made to visually demonstrate the differences in images generated by each method. Fig. 5 showcases images with the lowest SSIM, closest to the average SSIM, and the highest SSIM scores when using our method, compared to the outcomes when other methods are applied to the same images. We have also magnified some details to facilitate an intuitive comparison of the primary methods.
We also noted that the scores for the KITTI dataset are lower compared to those of the Middlebury dataset. However, if we convert the best scores into percentages relative to the Stable Diffusion reference scores, the results are as follows. For the Middlebury dataset, when SSIM = 0.551, it is 65.1% of the best score of 0.847, and for LPIPS = 0.173, the reference score of 0.046 constitutes 26.6% of the best score of 0.173 (the higher the percentage, the better). Similarly, for the KITTI dataset, SSIM is 63.1% of the reference score of 0.762, and the reference score for LPIPS of 0.072 is 35.1% of the best score. The model actually performs better on the KITTI dataset in terms of LPIPS. Another possible reason for this is the larger baseline distance $B$ of the cameras used to capture the KITTI dataset images, which in turn requires a larger scale factor $s$ (KITTI $s=20$ , Middlebury $s=9$). This larger scale factor means that, when generating stereo image pairs, the corresponding pixels in the KITTI dataset images have to move a greater distance, resulting in more extensive blank areas.
@@
Additionally, we compared the time consumption of different methods for generating a single stereo image pair on a GTX3090 GPU. The results of this comparison are in Table 2. Our method offers the capability to quickly generate high-quality stereo image pairs in a lightweight manner.
Table 2: Time cost for different methods in seconds. We measure the total time consumed for each usage scenario, including the time taken to generate the images using Stable Diffusion. The scenarios are text to stereo image (T2SI), depth to stereo image (D2SI), and image to stereo image (I2SI). The time in parenthesis is the cost excluding the time spent on generating images with Stable Diffusion. For D2SI, our method, being directly integrated into Stable Diffusion, requires only a single pass of sampling to generate stereo image pairs. In I2SI, our method requires use of null-text inversion [43] to obtain $x_{t}$, resulting in an extra 23 seconds of time expenditure.
 Methods T2SI D2SI I2SI
 3D Photography [12] 245 (231) 247 (231) 231
 Repaint[45] 338 (324) 340 (324) 324
 Ours 32 (18) 18 40 (17)
4.2 User evaluations
In our user tests, we adopted a more practical and user-centric approach. User input text prompts to generate stereo image pairs using Stable Diffusion. For benchmarking, we compared this with other methods by generating the left-side images using Stable Diffusion, obtaining the corresponding disparity maps via a depth estimation model, and then using the respective methods to generate stereo image pairs. We utilized Google Cardboard and presented the stereo images on mobile phones, inviting participants to assess the image quality and correctness of the 3D perception. Ratings ranged from 0 to 5, with 5 being the highest and 0 being the lowest. Some test pictures are shown in Fig. 12 of the supplemental document.
The results of the user tests showed that our method has the highest average but did not significantly outperform the others. This was anticipated, as when viewing stereo images, people tend to focus more on the overall image rather than the details. In terms of ease of use, our proposed method has a clear advantage. It is simpler, does not require an additional inpainting model, and can be seamlessly integrated with Stable Diffusion.
4.3 Ablation study
@@
Table 3: Ablation study on Middlebury and KITTI. In the ’Disparity Map’ column, ’GT’ and ’Pseudo’ respectively indicate the use of groundtruth disparity maps or disparity maps generated by a Depth Estimation Model. In the ’Technique Applied’ column, ’Attn Layer,’ ’SPSMD,’ and ’Deblur’ represent the use of Self-Attention Layers Modification, Symmetric Pixel Shift Masking Denoise, and Deblur techniques, respectively. The symbol ’✓’ denotes the adoption of these respective techniques. Bold numbers represent the best scores for that column. When employ Attn Layer and SPSMD together, LPIPS has a better score, but the effect of Deblur varies from image to image. When the LPIPS scores are comparable, the higher SSIM score indicates the better similarity, as an example shown in Fig. 7.
@@
GT Pseudo Attn Layers SPSMD Deblur PNSR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PNSR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
Figure 7: Ablation example of Middlebury (up) and KITTI (down). In the images, ’P’ and ’G’ respectively denote whether the image was guided by a Pseudo disparity map or a Groundtruth disparity map. ’A’, ’S’, and ’D’ indicate the use of Attention layers modification, Symmetric Pixel Shift Masking Denoise, and Deblur technique, respectively. The lower scores associated with the use of Groundtruth disparity maps in Middlebury may be attributed to their generally higher precision and complexity. This heightened detail can render pixel shift operations during image generation more intricate and sensitive. Our Stereo Pixel Shift operation is executed within a smaller latent space (64×64), where minor pixels, such as those around tree trunks and leaves, might be overlooked. In contrast, disparity maps generated by depth estimation models, with their lower precision, are more conducive to Pixel Shift in the latent space without sacrificing image detail.
@@
We conducted ablation studies on the proposed method to evaluate the impact of images guided by either Groundtruth disparity maps or Pseudo disparity maps (generated by a depth estimation model), as well as the effects of using Symmetric Pixel Shift Masking Denoise, Attention Layer Modification, and Deblur techniques on PSNR, SSIM, and LPIPS scores. The results are shown in Table 3. Fig. 7 presents a visual representation of an example from the Middlebury dataset and KITTI to intuitively demonstrate the impact of each factor on the image generation outcomes, explaining the reason that scores using Groundtruth disparity maps in the Middlebury dataset are unexpectedly lower than those using Pseudo disparity maps.

### v2

@@ 3.3 Self-Attention layers modification
As numerous studies have attempted to modify the attention mechanisms within Stable Diffusion to achieve the goal of modifying the original images [3, 10, 36, 22, 42], we tackle this challenge by utilizing both Unidirectional and Bidirectional Self-Attention mechanisms. This method eliminates the need for fine-tuning the model to adjust its weights. Refer to supplementary materials for detailed explanation.
@@
Our modified approach is listed in Algorithm 1. The term $\epsilon_{\theta}((z_{t},z^{\prime}_{t}),t,\mathcal{C})$ represents the computation of a single step $t$ of the diffusion process, which yields the noisy image $z_{t-1}$ and the attention map $M_{t}$. Here, $(z_{t},z^{\prime}_{t})$ denote the left and right latent variables, respectively. In our implementation, these latent variables are stacked together along the batch size dimension. However, they are presented separately here for ease of explanation. The expression $\epsilon_{\theta}((z_{t},{z^{\prime}_{t}}),t,\mathcal{C})\{M^{\prime}\leftarrow\widehat{M^{\prime}_{t}}\}$ denotes the diffusion step where the attention map $M$ is superseded by an additional map $\widehat{M}$. We define the function $\operatorname{Edit}\left(M_{t},M^{\prime}_{t},t\right)$ as a general edit function, designed to process the $t$-th attention maps of the left and right latent variables.
@@
Figure 4: The same image generated using different methods. The rows present, respectively, the image with the lowest (worst) SSIM score, the image closest to the average SSIM score, and the image with the highest (best) SSIM score generated using our method. The other methods are represented solely by their results on these specific images, which do not necessarily reflect the best, average, or worst SSIM scores achievable by those methods. We do this to facilitate a direct comparison of the effects of each method on the same image. We also provide LPIPS scores for reference and close-ups of the images generated by the primary benchmark methods for inspection of details.
@@
Table 1: Quantitative evaluation results for Middlebury and KITTI: the results of generating right-side images from left-side images and disparity maps using different methods. We assess the similarity between the generated and the original images using PSNR, SSIM, and LPIPS. ‘GT’ indicates the use of ground truth disparity maps, while ‘pseudo’ denotes the use of disparity maps generated by a depth estimation model. Scores presented in bold indicate the best performance. The numbers in the top right represent the best scores, while those in the bottom right indicate the worst scores.
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
3.4 Application scenarios
As shown in Figure 1, our method is compatible with various types of Stable Diffusion models, enabling it to: (a) produce the corresponding right-side image from an existing left-side image; (b) generate stereo images from text prompts; (c) produce the corresponding right-side image from an existing left-side image, where the pair shares the same composition but differs in content. For text-to-image and depth-to-image tasks, the initial noise is randomly generated. Thus, it is sufficient to apply a pixel shift to the denoised noise after a specific denoising step, as illustrated in Figure 2. However, for generating stereo image pairs of an existing image, it is necessary to use null-text inversion [19] to obtain the latent space of the original image. A straightforward inversion technique was proposed for DDIM sampling [5, 32]. This technique is grounded in the hypothesis that the ordinary differential equation (ODE) process is reversible, especially in scenarios involving small step sizes. The diffusion process is executed in reverse, meaning the transition is from $z_{0}$ to $z_{T}$, contrary to the typical $z_{T}$ to $z_{0}$ progression:
$z_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\,z_{t}$ (9)
$\quad{}+\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)\varepsilon_{\theta}\!\left(z_{t},t,\mathcal{C}\right).$
To address the inefficiency of mapping each noise vector to a single image, we start with a default DDIM inversion at $w=1$ as the pivot trajectory. Subsequently, we optimize around this trajectory using a standard guidance ratio of $w>1$. In practical applications, individual optimizations are conducted for each step $t$ during the diffusion process, aiming to closely approximate the initial trajectory $z^{*}$:
$\min\left\|z_{t-1}^{*}-z_{t-1}\right\|_{2}^{2}\,,$ (10)
where $z_{t-1}$ represents the intermediate result of the optimization. We thus substitute the default blank text embedding with an optimized embedding. This is because the generated results are significantly influenced by the unconditional prediction [19].
4 Experiments
We have compared our results with traditional methods such as ‘leave blank’ and ‘stretch’. Additionally, we have selected the 3D Photography techniques of Shih et al. [31] for comparison, as well as the RePaint method of Lugmayr et al. [16], which involves using Stable Diffusion for inpainting images processed by the traditional ‘leave blank’ method. It is important to emphasize that RePaint is not inherently designed for generating stereo image pairs. However, we believe that employing inpainting techniques to fill in the blank areas after creating stereo images is a very straightforward and common approach. Thus, we have chosen to compare with the latest model that achieves good results in various metrics within the same Stable Diffusion framework. This comparison is intended to demonstrate the innovation and advantage of our method.
4.1 Quantitative evaluation
@@
Since currently no metrics exists specifically for the comparison of stereo image pair generation, we quantitatively evaluate our results using the Middlebury [30] and KITTI [17] datasets. We evaluate the performance by generating the right-side image from the left-side image and its disparity map. We then compare the model-generated right-side image with the ground truth image. We calculated the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) between the generated image and the ground truth. The results are in Table 1 with some visuals in Figure 4. We provide the settings used for each method, more comparison figures, and a detailed explanation of the quantitative evaluation results in the supplemental document.
@@
In Table 3, we compare the time consumption of different methods for generating a single stereo image pair using an NVIDIA RTX3090 graphics card. Our method offers the capability to quickly generate high-quality stereo image pairs with direct integration into Stable Diffusion.
@@
Table 2: Ablation study on Middlebury and KITTI. In the ‘Disparity Map’ column, ‘GT’ and ‘Pseudo’ respectively indicate the use of groundtruth disparity maps or disparity maps generated by a Depth Estimation Model. In the ‘Technique Applied’ column, ‘Attn Layer,’ ‘SPSMD,’ and ‘Deblur’ represent the use of Self-Attention Layers Modification, Symmetric Pixel Shift Masking Denoise, and Deblur techniques, respectively. The symbol ‘✓’ denotes the adoption of these respective techniques. Bold numbers represent the best scores for that column. When employing Attn Layer and SPSMD together, LPIPS has a better score, but the effect of Deblur varies from image to image. When the LPIPS scores are comparable, the higher SSIM score indicates the better similarity. An example is shown in Figure 5.
@@
GT Pseudo Attn Layers SPSMD Deblur PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
Table 3: Time cost for different methods in seconds. We measure the total time consumed for each usage scenario, including the time taken to generate the images using Stable Diffusion. The scenarios are text to stereo image (T2SI), depth to stereo image (D2SI), and image to stereo image (I2SI). The time in parenthesis is the cost excluding the time spent on generating images with Stable Diffusion. For D2SI, our method, being directly integrated into Stable Diffusion, requires only a single pass of sampling to generate stereo image pairs. In I2SI, our method requires the use of null-text inversion [19] for $x_{t}$, resulting in an extra 23 seconds.
Methods T2SI D2SI I2SI
3D Photography [31] 245 (231) 247 (231) 231
Repaint[16] 338 (324) 340 (324) 324
Ours 32 (18) 18 40 (17)
@@
Figure 5: Ablation example of Middlebury (up) and KITTI (down). In the images, ‘P’ and ‘G’ respectively denote whether the image has been guided by a Pseudo disparity map or a Groundtruth disparity map. ‘A’, ‘S’, and ‘D’ indicate the use of Attention layers modification, Symmetric Pixel Shift Masking Denoise, and Deblur technique, respectively. The lower scores associated with the use of Groundtruth disparity maps in Middlebury may be attributed to their generally higher precision and complexity. This heightened detail can render pixel shift operations during image generation more intricate and sensitive. Our Stereo Pixel Shift operation is executed within a smaller latent space (64×64), where minor pixels, such as those around tree trunks and leaves, might be overlooked. In contrast, disparity maps generated by depth estimation models, with their lower precision, are more conducive to Pixel Shift in the latent space without sacrificing image detail.
4.2 User evaluations
In our user tests, we adopted a more practical and user-centric approach. The users’ input were text prompts used to generate stereo image pairs using Stable Diffusion. For benchmarking, we compared this with other methods by generating the left-side images using Stable Diffusion, obtaining the corresponding disparity maps via a depth estimation model, and then using the respective methods to generate stereo image pairs. We utilized Google Cardboard and presented the stereo images on mobile phones, inviting participants to assess the image quality and correctness of the 3D perception. Ratings ranged from 0 to 5, with 5 being the highest and 0 being the lowest. Some test pictures are shown in Figure 12 of the supplemental document.
The results of the user tests showed that our method has the highest average but it did not significantly outperform the others. This was anticipated, as when viewing stereo images, people tend to focus more on the overall image rather than the details. In terms of ease of use, our proposed method has a clear advantage. It is simpler, does not require an additional inpainting model, and can be seamlessly integrated with Stable Diffusion.
4.3 Ablation study
@@
We conducted ablation studies on the proposed method to evaluate the impact of images guided by either Groundtruth disparity maps or Pseudo disparity maps (generated by a depth estimation model), as well as the effects of using Symmetric Pixel Shift Masking Denoise, Attention Layer Modification, and Deblur techniques on PSNR, SSIM, and LPIPS scores. The results are shown in Table 2. Figure 5 presents a visual representation of an example from the Middlebury dataset and KITTI to intuitively demonstrate the impact of each factor on the image generation outcomes, explaining the reason that scores using Groundtruth disparity maps in the Middlebury dataset are unexpectedly lower than those using Pseudo disparity maps.
5 Limitations and Discussion
Our method relies on the disparity map. If the results generated by other depth estimation models are inaccurate, our method will also be unable to produce high-quality stereo images. Furthermore, when using high-precision disparity maps obtained from device measurements, the results may not be entirely satisfactory, as shown in Table 2 and Figure 5.
When using our depth to stereo image model, one may observe overlapping areas in the generated images. This issue might stem from the LatentDepth2ImageDiffusion model we used, which tends to fill blank areas with pixels from adjacent main subjects rather than background elements. In such cases, a better-quality image can be generated by first generating a single image using the Depth2Image model, and then applying our Image to Stereo Image Pairs method, as illustrated in Figure 7.
We found that our method can be used for inpainting tasks with the original text prompt to a Stable Diffusion image model. We conducted a simple test where, after obtaining $\bm{x}_{t}$ using null-text inversion, we applied various masking ratios to the right side and tested whether Stable Diffusion could fill in the blank areas within the mask during denoising. The results, as shown in Figure 8, indicate that our method is somewhat effective for inpainting when a smaller area of the image is masked. However, when a larger portion of the image is masked, the inpainting results exhibit a strong patchwork appearance. Applying our method to inpainting tasks might require further modifications to both the model and the technique.
6 Conclusion
Supplementary Material
7 Quantitative evaluation experiments setting
After rendering with these settings, we obtain the left and right images of the stereo image.
Given that Stable Diffusion can only generate images of $512\times 512$ resolution, and the Middlebury dataset images are about 5 million pixels, we scaled both the dataset images and the corresponding depth maps to $512\times 512$. For the Middlebury dataset, whose groundtruth disparity maps are noisy, we applied a Gaussian blur with a radius of 3 to smooth the disparity maps. Regarding the KITTI dataset, where the image size is $375\times 1242$ with an aspect ratio of approximately 3.3, directly scaling images to $512\times 512$ could lead to excessive stretching, negatively impacting many models’ performance. Therefore, we proportionally scaled the images to $512\times 1696$ and then applied a center crop to $512\times 512$. Because a null-text inversion technique is required, we used the Stable Diffusion version 1.5 for this test, setting the denoising steps to 50.
For the 3D photography method [31], we used the disparity map generated by the integrated MiDaS model [27] within its framework instead of the groundtruth disparity map. This was due to the extensive time required—up to two hours—for mesh reconstruction of a single image using the groundtruth disparity map with 3D photography. We hypothesize that this inefficiency arises when 3D photography attempts to reconstruct stereo image pairs from the disparity map, necessitating operations like breaking up discontinuous vertices in the mesh. Such processes become computationally intensive when the groundtruth disparity map is excessively noisy, leading to a proliferation of isolated vertices that consume substantial CPU resources. For the purpose of benchmarking and considering the rarity of obtaining groundtruth disparity maps in practical scenarios, we evaluated the results using both groundtruth disparity maps (denoted as GT disparity) and pseudo disparity maps generated by depth estimation models (denoted as Pseudo disparity). The depth estimation model we employed was DPT [26]. Since the use of Deblur results in lower scores, neither method employed deblur; details can be found in Sec. LABEL:sec:ablation. When creating stereo images using RePaint [16], we generate a mask for the blank areas left after moving the left-side image and then perform inpainting on the masked areas. The inet256 model was utilized for this purpose, and was trained on ImageNet. Since RePaint’s maximum supported output image size is $256\times 256$, we downsized the images to $256\times 256$ before conducting inpainting. However, considering that all other methods are evaluated at a $512\times 512$ resolution, for fairness, we only upscaled the inpainted area within the mask from $256\times 256$ to $512\times 512$, while maintaining the original resolution for the area outside the mask.
8 Analysis of Quantitative evaluation results
The use of the null-text inversion [19] technique inherently causes distortion in images. On the Middlebury dataset, reference scores (for images generated by Stable Diffusion to be the same as the input) are: PSNR = 27.967, SSIM = 0.847, LPIPS = 0.046. The reference scores for the KITTI dataset are: PSNR = 25.615, SSIM = 0.762, LPIPS = 0.072. These scores represent the best possible outcomes achievable with the method we proposed. The quantitative analysis results, as seen in Table 1, indicate that our proposed method achieves state-of-the-art scores on both the datasets. Furthermore, as illustrated in Fig. 9, we selected images representing the best LPIPS, those closest to the average LPIPS, and the worst LPIPS from each method. This selection was made to visually demonstrate the differences in images generated by each method. Fig. 4 showcases images with the lowest SSIM, closest to the average SSIM, and the highest SSIM scores when using our method. We also include a comparison to the outcomes when other methods are applied to the same images. We have also magnified some details to facilitate an intuitive comparison of the primary methods.
We also noted that the scores for the KITTI dataset are lower compared to those of the Middlebury dataset. However, if we convert the best scores into percentages relative to the Stable Diffusion reference scores, the results are as follows. For the Middlebury dataset, with SSIM = 0.551, it is 65.1% of the best score of 0.847, and for LPIPS = 0.173, the reference score of 0.046 constitutes 26.6% of the best score of 0.173 (the higher the percentage, the better). Similarly, for the KITTI dataset, SSIM is 63.1% of the reference score of 0.762, and the reference score for LPIPS of 0.072 is 35.1% of the best score. The model actually performs better on the KITTI dataset in terms of LPIPS. Another possible reason for this is the larger baseline distance $B$ of the cameras used to capture the KITTI dataset images, which in turn requires a larger scale factor $s$ (KITTI $s=20$ , Middlebury $s=9$). This larger scale factor means that, when generating stereo image pairs, the corresponding pixels in the KITTI dataset images have to move a greater distance, resulting in more extensive blank areas.
9 Analysis of Ablation